OpenStudy (anonymous):

A baseball is thrown up into the air. The ball's height in feet is modeled by the equation h(t) = -16t2 + 32t + 10. (h = height, t = time in seconds) Find the maximum height of the ball. a) 2.3 feet b) 23 feet c) 10 feet d) 26 feet

OpenStudy (anonymous):

You can differentiate it to find the maximum. \(h'(t)=( -16*2)t+32\)

OpenStudy (anonymous):

after differentiaton..... place the obtained t value in h(t)..... u get the maximum height!!!

OpenStudy (anonymous):

\[v(t)=h ^\prime(t)=0\] \[-32t+32=0\] \[t=1\] after 1 second the ball is at its maximum \[h(1)=-16(1)^{2}+32(1)+10=26\]

OpenStudy (anonymous):

at the maximum velocity is 0

OpenStudy (anonymous):

I Got The answer